# Augmented Lagrangian methods produce cutting-edge magnetic coils for stellarator fusion reactors

### v1

@@ III.3 Stellaris: a Stable Quasi-Isodynamic Stellarator (SQUID)
Here two alternative configurations are presented, one with 5 coils and the other with 6 coils per half-field period. In the former case, setting similar thresholds as the characteristics of Proxima’s coilset except for the intercoil distance, yields a configuration that improves on the magnetic field error average by 25%, as well as giving an additional 13 cm of space between coils which allows for better handling of eventual inaccuracies from manufacturing, assembly, thermal contraction or magnetic forces which appear 5% larger. The coilset is displayed on Figure 4. The main parameters are summarized on 4. Taking 200 turns per coil in the 6-coil solutions as reference and fixed total current per half-field period, it is possible to extrapolate the amount of HTS length required for the 2 solutions. While the two solutions come at a slightly higher cost of HTS tape, these configurations can enable extra access to the plasma, and facilitate reactor maintenance, which is a cost driver. Moreover, the 6 coil case improves on both the 5 coil configuration and Proxima’s coils with a field error that is cut by around half, linear forces on the HTS turns that are 29% lower, similar distances between coils and between the plasma surface and the conductor as Proxima. Note that this is however a conservative estimate of the amount of HTS required as it assumes that the tape is always operated at the limit of arbitrarily set engineering requirements for the critical current.
In order to achieve the reduced force configuration, the curvature and length curvatures constraints were relaxed, allowing the total length of the conductors to reach 145 m. The corresponding Boozer plots which serve as a verification for the departure from the target isodynamicity are shown in Appendix A.
Property | #1 | #2 | Proxima |
# of coils per hfp | 5 | 6 | 6 |
Max | |||
Max Force [MN/m] | |||
Total Length [m] | 130 | 145 | 138 |
Min CC Distance [m] | 0.8 | 0.7 | 0.67 |
Min CS Distance [m] | 1.38 | 1.3 | 1.37 |
Max [] | 1.57 | 1.6 | 1.57 |
Max MSC [] | 0.3 | 0.35 | 0.34 |
HTS Length [km] | 250 |
As shown in Table 4, the Stellaris configuration is improved in essentially all aspects by the two different coil solutions found here, except potentially for the amount of HTS required. Nonetheless, final decisions about coil and HTS stack design require full finite-element modeling of the coils and structural supports beyond the scope of this work.
@@ III.4 Wendelstein 7-X (W7-X)
As is illustrated in Table 5, all the configurations are optimized to possess lower curvatures than the original W7-X coilset. The results are visible in the coils shown in Figure 5, as all the newly optimized coils appear smoother and with less local curvature than the W7-X ones. Note that all configurations possess comparable lengths to the W7-X coils. This freedom is purposely imposed for this study, as the goal is to explore options to the current W7-X reactor coils and not to replicate it.
@@
The degree of omnigeneity for all coil sets displayed by the magnetic field strength contours is plotted in Appendix A. All W7-X alternatives show contours following the target field. As shown in Figure 5, the #2 and #3 alternatives possess four visibly simple coils out of five with only one coil type standing out as longer. It also appears that lower overall forces can be achieved by relaxing the length constraint on an individual coil type. To our current knowledge and given the current parameters of the W7-X coils displayed in Table 5, any of these three alternatives of W7-X could be considered as candidates for a W7-X-style research reactor.
@@ III.5 Helically Symmetric eXperiment (HSX)
The Helically Symmetric eXperiment (HSX) is a stellarator located at the University of Wisconsin in Madison. The device began operation in 1999, and was and still is the only device in the world with a quasi-helical stellarator field. It is a four field period stellarator with 48 copper coils, around 1T on axis and major radius of 1.4m. HSX was designed to demonstrate reduced direct orbit losses, Mercier stability, and reduced neoclassical transport. Due to its rotational transform of 1.05 on axis and quasi-helical field, its neoclassical confinement is similar to a tokamak [6]. It has also managed to demonstrate reduced flow damping and reduced electron heat transport [63, 64]. Unlike the QH stellarator equilibrium shown in Section III.2 which was optimized for low quasisymmetric error throughout the volume, HSX was only optimized for quasisymmetry on the last closed flux surface because at the time no robust optimization methods had been developed yet. Similarly to W7-X, HSX coils were obtained using a winding surface method, which yielded six coils per half field period with high curvatures. From the twisted filaments, a copper winding pack was constructed for the final reactor, these are shown in Figure 6. The procedure in this section does not aim at improving the forces, since the force requirements become much more stringent at reactor-scale. Given the smaller dimensions and relatively low magnetic fields of this device, we focus instead on finding simpler coils that are easier to manufacture. We propose three different configurations with four, five, and six coils per half-field period, summarized in Table 6. As expected, the average field error progressively increases with decreasing number of coils. This is consistent with the results from the previous QH, Stellaris and W7-X configurations.
@@
The three proposed configurations are mostly characterized by their reduction in curvature, which can be understood as a proxy for coil complexity. The reduction in mean squared curvature reaches 45% in configuration #1, and absolute normal curvature is decreased by 35% in configuration #3. Despite configuration #3 exhibiting the same number of coils per hfp, it manages to improve on the field accuracy while having simpler coils, the same clearance inbetween coils, and losing only 3 cm in the coil-to-surface distance. The forces are generally increased in all the optimized configurations, as they are not a target. It remains the sole aspect where the original HSX design is better.
Configuration #1 is an exemplary solution. Configuration #1 has reduced by two the number of coils per half field period. This level of coil reduction has not been tried in previous sections, as it is often expected that field accuracy will not be precise enough with so few coils per hfp. Amazingly, and despite the fewer coils per hfp, configuration #1 has reasonable field accuracy, and, compared with the original six-coil design, improved minimum coil-coil and coil-surface distances, lower curvatures, and achieves all of these improvements with 25% less total coil length. Despite the factor of two increase in field error compared to the original design, the magnetic field contours plotted in 11 still display comparable accuracy to the other candidates. As previously mentioned, less coils are of particular interest in a research scale reactors where additional diagnostics are usually designed throughout the operation time and need space for new ports. The additional space in between coils on the outboard side is noticeable on Figure 6. Less coils also means less machining time and workforce required in the case of HSX as the coils are made out of copper and usually require the operation of CNC-milling or other heavy machinery devices. This can directly translate to reduced costs, which is often a limiting factor in university scale reactors and more generally in any type of reactor.
@@ Acknowledgements.
This work was supported through grants from the Simons Foundation under award 560651 and from the Helmholtz Association Young Investigators Group program as project VH-NG-1430. We thank Jorrit Lion, Gabriel Plunk, Alan Goodman, Matt Landreman, and Benedikt Geiger for sharing plasma boundary configuration files, and Rory Conlin, Paul Huslage, David Bindel and Tom Simpson for useful discussions.Appendix A Equilibria verification

### v2

@@ III.3 Stellaris: a Stable Quasi-Isodynamic Stellarator (SQUID)
Here three alternative configurations are presented, one with 5 coils, two with 6 coils per half-field period. In the former case, setting similar thresholds as the characteristics of Proxima’s coilset except for the intercoil distance, yields a configuration that improves on the magnetic field error average by 25%, as well as giving an additional 13cm of space between coils which allows for better handling of eventual inaccuracies from manufacturing, assembly, thermal contraction or magnetic forces which appear 5% larger. The coilset is displayed on Figure 5. The main parameters are summarized on 4. Taking 200 turns per coil in the 6-coil solutions as a reference, and fixed total current per half-field period, it is possible to extrapolate the amount of HTS length required for the 2 solutions. While the two solutions come at a slightly higher cost of HTS tape, these configurations can enable extra access to the plasma, and facilitate reactor maintenance, which is a cost driver. Moreover, the 6 coil case #2 improves on both the 5 coil configuration and Proxima’s coils with a field error that is cut by around half, linear forces on the HTS turns that are 29% lower, similar distances between coils and between the plasma surface and the conductor as Proxima. Note that this is however a conservative estimate of the amount of HTS required as it assumes that the tape is always operated at the limit of arbitrarily set engineering requirements for the critical current. In order to achieve the reduced force configuration #2, the curvature and length curvatures constraints were relaxed, allowing the total length of the conductors to reach 145m. The corresponding Boozer plots which serve as a verification for the departure from the target isodynamicity are shown in Appendix B.
Furthermore, configuration #3 takes advantage of the new threshold formulation for the quadratic flux by relaxing it while making every other constraint stricter. This yields a configuration with field errors four times higher than Stellaris but improves essentially every engineering metric. This is of special interest since it is unclear if achieving accuracies of is at all possible at a reactor scale without auxiliary coils. Therefore, we choose to optimize for a coilset with higher errors but with for example more 32cm of spacing between coils leaving space for error correction coils. This is achieved while having simpler geometries with a 42% decrease in local curvature, a 38% decrease in mean squared curvature and 8m/turn shorter coils giving a total 6% decrease in HTS cost. Similarly to configuration #2, forces are also reduced by 29%.
| Property | #1 | #2 | #3 | Proxima |
| # of coils per hfp | 5 | 6 | 6 | 6 |
| 18.5 | ||||
| Max | 12.2 | |||
| Max Force [MN/m] | 0.79 | 0.64 | 0.64 | 0.9 |
| Total Length [m] | 130 | 145 | 130 | 138 |
| Min CC Distance [m] | 0.8 | 0.7 | 0.99 | 0.67 |
| Min CS Distance [m] | 1.38 | 1.3 | 1.39 | 1.37 |
| Max [] | 1.57 | 1.6 | 0.91 | 1.57 |
| Max MSC [] | 0.3 | 0.35 | 0.21 | 0.34 |
| HTS Length [km] | 250 | 232 | 208 | 221 |
As shown in Table 4, the Stellaris configuration is improved in essentially all aspects by the 3 different coil solutions found here. Notably, configuration #3 is able to improve all the engineering metrics, including 29% less forces and a 6% reduction in estimated HTS cost. Nonetheless, final decisions about coil and HTS stack design require full finite-element modeling of the coils and structural supports beyond the scope of this work.
@@ III.4 Wendelstein 7-X (W7-X)
As is illustrated in Table 5, all the configurations are optimized to possess lower curvatures than the original W7-X coilset. Here, the last closed flux surface of W7-X that does not contain coil ripple is used as target. The results are visible in the coils shown in Figure 6, as all the newly optimized coils appear smoother and with less local curvature than the W7-X ones. Note that all configurations possess comparable lengths to the W7-X coils. This freedom is purposely imposed for this study, as the goal is to explore options to the current W7-X reactor coils and not to replicate it.
@@
The degree of omnigeneity for all coil sets displayed by the magnetic field strength contours is plotted in Appendix B. All W7-X alternatives show contours following the target field. As shown in Figure 6, the #2 and #3 alternatives possess four visibly simple coils out of five with only one coil type standing out as longer. It also appears that lower overall forces can be achieved by relaxing the length constraint on an individual coil type. To our current knowledge and given the current parameters of the W7-X coils displayed in Table 5, any of these three alternatives of W7-X could be considered as candidates for a W7-X-style research reactor.
@@ III.5 Helically Symmetric eXperiment (HSX)
The Helically Symmetric eXperiment (HSX) is a stellarator located at the University of Wisconsin in Madison. The device began operation in 1999, and was and still is the only device in the world with a quasi-helical stellarator field. It is a four field period stellarator with 48 copper coils, around 1T on axis and major radius of 1.4m. HSX was designed to demonstrate reduced direct orbit losses, Mercier stability, and reduced neoclassical transport. Due to its rotational transform of 1.05 on axis and quasi-helical field, its neoclassical confinement is similar to a tokamak [6]. It has also managed to demonstrate reduced flow damping and reduced electron heat transport [63, 64]. Unlike the QH stellarator equilibrium shown in Section III.2 which was optimized for low quasisymmetric error throughout the volume, HSX was only optimized for quasisymmetry on the last closed flux surface because at the time no robust optimization methods had been developed yet. Similarly to W7-X, HSX coils were obtained using a winding surface method, which yielded six coils per half field period with high curvatures. From the twisted filaments, a copper winding pack was constructed for the final reactor, these are shown in Figure 7. The procedure in this section does not aim at improving the forces, since the force requirements become much more stringent at reactor-scale. Given the smaller dimensions and relatively low magnetic fields of this device, we focus instead on finding simpler coils that are easier to manufacture. We propose three different configurations with four, five, and six coils per half-field period, summarized in Table 6. Note that the target configuration used in this section is a free-boundary surface, meaning that it already has coil ripple embedded into it, making it harder to optimize for different number of coils per hfp. The reference coilset used here is obtained by taking the center spline of finite-built coils, which accounts for the field error being non-zero. Nevertheless, we report comparably low field errors, and as expected, the average field error progressively increases with decreasing number of coils. This is consistent with the results from the previous QH, Stellaris and W7-X configurations.
@@
The three proposed configurations are mostly characterized by their reduction in curvature, which can be understood as a proxy for coil complexity. The reduction in mean squared curvature reaches 45% in configuration #1, and curvature is decreased by 35% in configuration #3. Despite configuration #3 exhibiting the same number of coils per hfp, it manages to improve on the field accuracy while having simpler coils, the same clearance inbetween coils, and losing only 3 cm in the coil-to-surface distance. The forces are generally increased in all the optimized configurations, as they are not a target. It remains the sole aspect where the original HSX design is better.
Configuration #1 is an exemplary solution. Configuration #1 has reduced by two the number of coils per half field period. This level of coil reduction has not been tried in previous sections, as it is often expected that field accuracy will not be precise enough with so few coils per hfp. Amazingly, and despite the fewer coils per hfp, configuration #1 has reasonable field accuracy, and, compared with the original six-coil design, improved minimum coil-coil and coil-surface distances, lower curvatures, and achieves all of these improvements with 25% less total coil length. Despite the factor of two increase in field error compared to the original design, the magnetic field contours plotted in 12 still display comparable accuracy to the other candidates. As previously mentioned, less coils are of particular interest in a research scale reactors where additional diagnostics are usually designed throughout the operation time and need space for new ports. The additional space in between coils on the outboard side is noticeable on Figure 7. Less coils also means less machining time and workforce required in the case of HSX as the coils are made out of copper and usually require the operation of CNC-milling or other heavy machinery devices. This can directly translate to reduced costs, which is often a limiting factor in university scale reactors and more generally in any type of reactor.
@@ Acknowledgements.
This work was supported through grants from the Simons Foundation under award 560651 and from the Helmholtz Association Young Investigators Group program as project VH-NG-1430. We thank Jorrit Lion, Gabriel Plunk, Alan Goodman, Matt Landreman, and Benedikt Geiger for sharing plasma boundary configuration files, and Rory Conlin, Paul Huslage, David Bindel and Tom Simpson for useful discussions.Appendix A Quasisymmetry Metric
In this work, the metric used to quantify the departure from quasisymmetry is the two-term metric [9, 65], which does not require a transformation into Boozer coordinates and can be expressed as
Here it appears as a weighted sum across all specified flux surfaces , with the weights by default set to 1, is the magnetic field amplitude, are parameters specifying the class of quasisymmetry (either QH or QA), is the toroidal flux function, and is the poloidal and the toroidal currents inside the surface.
Appendix B Equilibria verification